### v2

@@ Abstract
When nanoparticles (NPs) are introduced into a biological solution, layers of biomolecules form on their surface, creating a corona. Understanding how the protein’s structure evolves into the corona is essential for evaluating the safety and toxicity of nanotechnology. However, the influence of NP properties on protein conformation is not well understood. In this study, we propose a new method that addresses this issue by analyzing multi-component spectral data using machine learning (ML). We apply the method to fibrinogen, a crucial protein in human blood plasma, at physiological concentrations while interacting with hydrophobic carbon or hydrophilic silicon dioxide NPs, revealing striking differences in the temperature dependence of the protein structure between the two cases. Our unsupervised ML method a) does not suffer from the challenges associated with the curse of dimensionality, and b) simultaneously handles spectral data from various sources. The method offers a quantitative analysis of protein structural changes upon adsorption and enhances the understanding of the correlation between protein structure and NP interactions, which could support the development of nanomedical tools to treat various conditions.
@@ 1 Introduction
Within the main constituents of the corona formed when NPs are exposed to human plasma, we find Fibrinogen (Fib) [5], one of the most abundant glycoproteins in human blood plasma with a concentration of about 1.5–4 g/L.[6]. Synthesized in the liver by hepatocytes, it is essential for the coagulation cascade and also plays a significant role in blood viscosity, blood flow, and various other biological functions [7]. During the last decades, many studies have focused on the correlation of unusually elevated levels of Fib with an increased risk of experiencing health disorders, such as cardiovascular diseases [6].
@@
Here, by combining different experimental techniques with Machine Learning (ML) analysis, we characterize the structural changes in Fib over temperature when exposed to hydrophobic carbon NPs (CNP) and hydrophilic silicon dioxide (SiO2) NPs. We employ several spectroscopic techniques: UltraViolet Resonance Raman (UVRR) spectroscopy [19], Circular Dichroism (CD) [20], and UV absorption [21] spectroscopy.
@@
All these techniques are non-destructive and do not require chemical labeling. For instance, by setting the excitation wavelength in UVRR to the deep UV region (226 nm), the Raman signals associated with aromatic amino acids, such as tryptophan (Trp) and tyrosine (Tyr), are selectively enhanced, providing insights into the protein’s tertiary structure and how its chromophores interact with their surroundings [22, 23]. On the other hand, CD spectroscopy, which is based on the differential absorption of left- and right-hand circularly polarized light, can indicate the presence of disorder in the protein’s secondary structure [24, 25]. Lastly, UV absorption spectra are very sensitive to the presence of aggregates and precipitates, which can elastically scatter light according to Rayleigh’s law [19].
@@
We analyze our high-dimensional, multi-component data from various spectroscopic techniques using a novel unsupervised ML workflow designed to quantify and visualize how protein conformational changes occur as temperature varies. The novelty of our proposed ML model rests on two key algorithms: the Wasserstein distance [26, 27] and t-distributed stochastic neighbor embedding (t-SNE) [28, 29, 30].
@@
These statistical tools are particularly advantageous because both remain unaffected by the curse of dimensionality, i.e., the challenges and limitations of working with high-dimensional data, such as overfitting, sparse distribution, computational complexity, and saturation of distances [31, 32, 33]. Indeed, high-dimensional data render standard similarity tools, such as the Euclidean and Manhattan distances commonly used in cheminformatics [34], ineffective.
@@
Furthermore, t-SNE successfully replaces principal component analysis (PCA) [35, 36, 31], which is routinely employed to visualize spectral data in two-dimensional space. PCA cannot effectively handle data from diverse sources, mainly because, by preserving large distances arising from different spectra, it misses interesting structures, such as clustering, at smaller distances [^1] This is why PCA is susceptible to outliers in the data. In contrast, t-SNE overcomes this issue. When PCA is used as an initialization step, t-SNE can preserve both global and local data structures, yielding a meaningful embedding [30, 37].
@@
Our multi-component data are organized into vectors with 856 dimensions, measured between 22 and 86C at temperature intervals of $\Delta T=4^{\circ}$. Overall, our analysis reveals that the protein structural disorder gradually increases for Fib adsorbed onto CNPs as the temperature $T$ rises, suggesting a ’frozen’ partially unfolded structure that is only weakly dependent on $T$. In contrast, the structural behavior for Fib adsorbed onto SiO2NPs is similar to that of free Fib, both characterized by a significant increase in disorder around 50C. These findings are further supported by the t-SNE cluster analysis. For free Fib, the substantial increase in disorder is experimentally understood as a direct result of Fib aggregation. Therefore, our novel ML analysis of multi-component spectral data offers an experimental confirmation of theoretical predictions, providing insights into the interaction mechanisms between proteins and NPs. It highlights the importance of NP surface properties in protein adsorption and stability, and represents a new and solid method to analyze the controversial structural changes during biomolecular corona evolution.
@@
Figure 2: Physicochemical characterization of the NPs and the Fib corona. (a) Transmission Electron Microscopy (TEM) image of CNPs with a nominal diameter of $\sim$120 nm. After image processing, they were calculated to have a core diameter of (118.3 $\pm$ 40.8) nm. (b) DLS distribution for pristine (’plain’) CNPs and in solution with Fib at WR between 0.1 and 1.0. In all cases, the polydispersity index (PDI) is less than 0.12. (c) DCS analysis for the corresponding cases. (d) TEM image of SiO2NPs with a nominal diameter of 100 nm. Their calculated size of (95.4 $\pm$ 7.8) nm. (e) DLS of SiO2NPs with PDI of 0.01, and (f) the corresponding DCS analysis.
@@ 2.1 Experimental results
We follow the experimental protocols developed in our lab to obtain the protein corona complexes formed on top of carbon and silica NPs [38]. We first characterize the pristine NPs by Transmission Electron Microscope (TEM). By using an electron beam to image samples, TEM is able to provide high-resolution images and to characterize the size populations of NPs (Fig.2a and d). Furthermore, by Dynamic Light Scattering (DLS), we characterize their polydispersity index (PDI). In DLS, a laser tracks the NPs’ movement, due to Brownian diffusion, allowing the measurement of the autocorrelation function and diffusion coefficient, which in turn determine the particles’ hydrodynamic size and PDI. We find that both CNPs and SiO2NPs are monodisperse in aqueous solution and stable against aggregation, even though the former are hydrophobic and the latter are hydrophilic (Fig.2a and d).
Next, we check whether the solutions of NPs covered by a corona of Fib are stable at room temperature. Previous experiments have shown that the addition of proteins in solution can induce NP aggregation [39], and that aggregation in simple media can occur at much lower protein concentrations than in complex media with many components [40].
Therefore, we analyze the stability of the NP-protein corona complexes against aggregation after incubation at different protein/NP weight ratios (WR from 0.1 to 1.0) using various techniques. By employing DLS and Differential Centrifugal Sedimentation (DCS), we do not observe aggregates for CNPs at the tested WRs (Fig.2 b and c). However, for the same range of WRs, aggregation systematically occurs with SiO2NP (Fig.2 e and f).
Therefore, the difference in the NPs’ chemistry likely impacts differently the protein adsorption and the aggregative propensity. Next, we set our experimental conditions to WR=1 and compare the cases of both CNP (colloidally stable) and SiO2NP (colloidally unstable) with a free Fib solution at various temperatures.
@@
Next, our DLS analysis indicates that the hydrodynamic diameter of bulk Fib and Fib with silica NPs varies with temperature. In both cases, there is a heterogeneous PDI, with aggregation occurring at 52C and above.
In contrast, the sample containing Fib and carbon NPs shows a homogeneous PDI at any temperature. However, at 52C, it exhibits a shift in PDI that may be related to a gradual denaturation of Fib, which affects the solution’s colloidal stability.
@@ 2.1.2 UV Resonance Raman Spectroscopy
UVRR spectroscopy is a powerful tool for detecting structural changes in organic biomolecules, such as proteins and peptides, without the need for any chemical label [42]. Compared to spontaneous Raman spectroscopy using visible excitation sources, the UVRR technique offers undeniable advantages. The vibrational profiles of proteins typically result in a complex superposition of several signals, making analysis difficult. Thanks to the selective resonance enhancement achieved by appropriately tuning the excitation wavelength in the deep UV range (200-300 nm), UVRR spectra of proteins can be simplified, and the signals arising from selected chromophores or molecular portions of the molecule can be more easily disentangled [43, 44, 45]. In this regard, using Synchrotron Radiation (SR) as an excitation source for UVRR experiments provides the added advantage of fine-tuning the excitation wavelength to select the best working conditions for the system under investigation [46, 47, 48]. Additionally, the high sensitivity of the UVRR technique is crucial for investigating proteins in more diluted conditions compared to visible Raman spectroscopy while maintaining satisfactory quality in the spectra. This, in turn, provides the opportunity to simultaneously extract information from the UVRR spectra of proteins in aqueous solutions, both from the Raman signals attributed to the polypeptide solute and the water molecules in the hydration shell [49, 47].
@@
We collect the UVRR spectra for Fib in PBS buffer, both in the presence and absence of NPs, using 226 nm as the excitation source (Fig. 4 a-d). This excitation energy falls within the absorption range of the aromatic rings of Tyrosine (Tyr) and Tryptophan (Trp) residues, resulting in UVRR spectra of Fib that are predominantly characterized by the vibrational signals of these aromatic amino acids [43, 50] (Fig. 4 a).
@@
In the Raman spectrum of the Fib protein, we can identify several key resonance Raman features associated with the Trp vibrational modes, including signals at 1012 $\mathrm{cm^{-1}}$ (Trp W16), which are attributed to the symmetric benzene/pyrrole out-of-phase breathing, at 1340-1360 $\mathrm{cm^{-1}}$ (Trp W7), which is attributed to the Fermi resonance between N1-C8 stretching in the pyrrole ring and combination bands of out-of-plane bending, and at 1554 $\mathrm{cm^{-1}}$ (Trp W3), which relates to the C2-C3 stretching mode of the pyrrole ring [51]. In the same spectral range of 1000-1800 $\mathrm{cm^{-1}}$, Tyr Raman bands are also detected, specifically the signals at 1176 $\mathrm{cm^{-1}}$ (Tyr Y9a), which are associated with in-plane C-H bending, and at 1616 $\mathrm{cm^{-1}}$ (Tyr 8a), which is due to the in-plane ring stretching of Tyr. Additionally, a signal at approximately 1770 $\mathrm{cm^{-1}}$ is observed in the UVRR spectrum of Fib, likely corresponding to a C=O stretching vibration.
@@
Figure 4: 226 nm-excited UVRR spectra collected for Fib 2mg/ml in PBS pH 7.4 (a), Fib with CNPs (b) and Fib with SiO2NPs (c) as a function of temperature, reported in the two spectral ranges 1000-1800 and 2800-3800 $\mathrm{cm^{-1}}$. The main features in the spectrum of Fib are labelled in the panel (a). Temperature range and steps are indicated in the colour scale at the right. Dashed lines in panels (b) and (c) represent the spectra of Fib alone at the lowest temperature of 22C.
@@
Since the aromatic rings of Trp and Tyr can participate in hydrophobic, cation-$\mathrm{\pi}$, and hydrogen-bond interactions with nearby residues or surrounding solvent molecules [51], the spectral feature parameters of the Raman bands associated with these residues provide valuable insights into their local environment. For instance, it has been reported that the vibrations of aromatic ring side chains in proteins can serve as sensitive markers for changes occurring in the tertiary structure of the proteins during unfolding and aggregation [23, 22, 52, 53, 54].
@@
It is noteworthy that, in the spectral range of 1000-1800 $\mathrm{cm^{-1}}$ (Fig. 4 a), the Raman signal from the water solvent is practically negligible compared to the vibrational signals attributed to the Fib protein. Conversely, the high frequency range between 2800 and 3800 $\mathrm{cm^{-1}}$ in the UVRR spectrum of Fib is entirely dominated by the intense spectral distribution characteristic of the OH stretching profile of water.
@@
Figure 5: Integrated area of the UVRR spectra in Fig. 4. (a) OH band area for Fib in PBS (black circles) and with CNPs (red pentagons), as well as SiO2NPs (green squares). (b)-(e) Areas of the peaks in the frequency intervals reported in Table 1 for Fib in PBS and with CNPs and SiO2NPs. In each panel, the frequency range is indicated as a label, and the symbols are the same as in panel (a). Where indicated, the experimental data have been fitted with a sigmoidal curve (full black lines).
The integrated area of the OH stretching band for Fib alone (Fig. 5 a) exhibits a clear sigmoidal trend as a function of temperature, indicating a transition of the protein occurring at approximately 57C. The sudden increase in this Raman signal above 50C likely marks the unfolding and aggregation of the protein driven by the temperature rise. The sigmoidal trend suggests that protein denaturation begins at around 50C and rapidly increases until 70C, at which point the intensity of the OH band area stabilizes. Consequently, aggregates may be forming within this temperature range.
When CNPs are added to Fib’s solution, the OH profile significantly decreases in intensity but retains its shape (Fig. 4 b). This may suggest that water is not able to penetrate the protein corona, likely due to protein unfolding on the surface of the NP. Additionally, the corresponding OH-band integrated area shows little to no dependence on $T$ (Fig. 5 a). This finding suggests that the Fib corona is so strongly adsorbed that even with a significant increase in the energy of the solution due to heating, the proteins remain bound to the CNP. Furthermore, the analysis provides no evidence of aggregation.
When SiO2NPs are in solution with Fib, we observe that the OH-band integrated area is intermediate between the two previous cases, with a $T$-dependence that is weaker than for free Fib (Fig. 5 a). Furthermore, the increase in the area occurs at a temperature close to that of bulk Fib. This behavior suggests that Fib absorbed onto SiO2NPs undergoes unfolding and aggregation similar to free Fib. Nevertheless, the analysis does not allow us to quantify the amount of unfolding and aggregation relative to the bulk case.
We also analyze the temperature dependence of the integrated area of some characteristic Raman signals in the fingerprint region 1000-1800 $\mathrm{cm^{-1}}$ assigned to Fib (Fig. 5 c-f). In particular, the unfolding and aggregation of proteins should alter the local packing of side chains and the solvent exposure of the residues, affecting the surrounding environment of Trp amino acids and the resonance Raman cross-section of the associated bands [50, 55, 56]. Therefore, by comparing their relative intensity for Fib in the absence and presence of NPs, we can qualitatively assess the effect of NPs on the $T$-dependence of the bands (Fig. 5 c-f).

### v3

@@ Abstract
When nanoparticles (NPs) are introduced into a biological solution, layers of biomolecules form on their surface, creating a corona. Understanding how the protein’s structure evolves into the corona is essential for evaluating the safety and toxicity of nanotechnology. However, the influence of NP properties on protein conformation is not well understood. In this study, we propose a new method that addresses this issue by analyzing multi-component spectral data using machine learning (ML). We apply the method to fibrinogen, a crucial protein in human blood plasma, at physiological concentrations while interacting with hydrophobic carbon or hydrophilic silicon dioxide NPs, revealing striking differences in the temperature dependence of the protein structure between the two cases. Our unsupervised ML method a) does not suffer from the challenges associated with the curse of dimensionality, and b) simultaneously handles spectral data from various sources. The method offers a quantitative analysis of protein structural changes upon adsorption. It enhances the understanding of the correlation between protein structure and NP interactions, which could support the development of nanomedical tools to treat various conditions.
@@ 1 Introduction
Within the main constituents of the corona formed when NPs are exposed to human plasma, we find Fibrinogen (Fib) [5], one of the most abundant glycoproteins in human blood plasma with a concentration of about 1.5–4 g/L.[6]. Synthesized in the liver by hepatocytes, it is essential for the coagulation cascade and also plays a significant role in blood viscosity, blood flow, and various other biological functions [7]. During the last decades, many studies have focused on the correlation of unusually elevated levels of Fib with an increased risk of experiencing health disorders, such as cardiovascular diseases [6]. Furthermore, Fib was shown to mediate adverse responses when interacting with biomaterials [8] or nanobiomaterials [9].
@@
Here, by combining different experimental techniques with Machine Learning (ML) analysis, we characterize the structural changes in Fib over temperature when exposed to carbon NPs (CNP) and silicon dioxide (SiO2) NPs. We use monodisperse nanoparticles of similar size but differing surface chemistry. CNPs and SiO2NP have a comparable geometrical and hydrodynamic diameter to prevent any effects caused by a differing surface curvature. Both exhibit a negative zeta potential at neutral pH [5, 21], indicating the presence of weakly acidic groups on the surface. In the case of CNPs, these species include acidic carboxylic or phenolic groups formed during synthesis, while for SiO2NP, they consist of surface hydroxyl groups (silanols, Si-OH). The density of acidic groups is expected to be in the same range for both samples (around 4-6 groups per nm2) [5]. However, the two surfaces exhibit significant differences. While in CNPs the oxygenated moieties are bound to a hydrophobic surface composed of C-C bonds, in the case of silica, silanols are attached to a more hydrophilic framework made up of polar Si-O-Si bridges. Therefore, while SiO2NPs are hydrophilic, CNPs are more hydrophobic.
We employ several spectroscopic techniques: UltraViolet Resonance Raman (UVRR) spectroscopy [22], Circular Dichroism (CD) [23], and UV absorption [24] spectroscopy.
@@
All these techniques are non-destructive and do not require chemical labeling. For instance, by setting the excitation wavelength to correspond with the electronic transition energy of the aromatic amino acids around 220-230 nm, the Raman signals associated with tryptophan (Trp) and tyrosine (Tyr) are selectively enhanced, providing insights into the protein’s tertiary structure and how its chromophores interact with their surroundings [25, 26]. On the other hand, CD spectroscopy, which is based on the differential absorption of left- and right-hand circularly polarized light, can indicate the presence of disorder in the protein’s secondary structure [27, 28]. Lastly, UV absorption spectra are very sensitive to the presence of aggregates and precipitates, which can elastically scatter light according to Rayleigh’s law [22].
@@
We analyze our high-dimensional, multi-component data from various spectroscopic techniques using a novel unsupervised ML workflow designed to quantify and visualize how protein conformational changes occur as temperature varies. The novelty of our proposed ML model rests on two key algorithms: the noisy-resistant Wasserstein distance [29, 30] and t-distributed stochastic neighbor embedding (t-SNE) [31, 32, 33].
@@
These statistical tools are particularly advantageous because both remain unaffected by the curse of dimensionality, i.e., the challenges and limitations of working with high-dimensional data, such as overfitting, sparse distribution, computational complexity, and saturation of distances [34, 35, 36]. Indeed, high-dimensional data render standard similarity tools, such as the Euclidean and Manhattan distances commonly used in cheminformatics [37], ineffective (Section 1 in Electronic Supplementary Information, ESI).
@@
Furthermore, t-SNE successfully replaces principal component analysis (PCA) [38, 39, 34], which is routinely employed to visualize spectral data in two-dimensional space. PCA cannot effectively handle data from diverse sources, mainly because, by preserving large distances arising from different spectra, it misses interesting structures, such as clustering, at smaller distances [^1] This is why PCA is susceptible to outliers in the data (Section 2 of ESI). In contrast, t-SNE overcomes this issue. When PCA is used as an initialization step, t-SNE can preserve both global and local data structures, yielding a meaningful embedding [33, 40].
@@
Our multi-component data are organized into vectors with 856 dimensions, measured between 22 and 86C at temperature intervals of $\Delta T=4^{\circ}$. Overall, our analysis reveals that the protein structural disorder gradually increases for Fib adsorbed onto CNPs as the temperature $T$ rises, suggesting a ’frozen’ partially unfolded structure that is only weakly dependent on $T$. In contrast, the structural behavior for Fib adsorbed onto SiO2NPs is similar to that of free Fib, both characterized by a significant increase in disorder around 50C. The t-SNE cluster analysis further supports these findings. For free Fib, the substantial increase in disorder is experimentally understood as a direct result of Fib aggregation. Therefore, our novel ML analysis of multi-component spectral data offers an experimental confirmation of theoretical predictions, providing insights into the interaction mechanisms between proteins and NPs. It highlights the importance of NP surface properties in protein adsorption and stability, and represents a new and solid method to analyze the controversial structural changes during biomolecular corona evolution.
@@
Figure 2: Physicochemical characterization of the NPs and the Fib corona. (a) Transmission Electron Microscopy (TEM) image of CNPs with a nominal diameter of $\sim$120 nm. After image processing, they were calculated to have a core diameter of (118.3 $\pm$ 40.8) nm. (b) DLS distribution for pristine (‘plain’) CNPs and in solution with Fib at WR between 0.1 and 1.0. In all cases, the polydispersity index (PDI) is less than 0.12. (c) DCS analysis for the corresponding cases. (d) TEM image of SiO2NPs with a nominal diameter of 100 nm. Their calculated size of (95.4 $\pm$ 7.8) nm. (e) DLS of SiO2NPs with PDI of 0.01, and (f) The corresponding DCS analysis.
@@ 2.1 Experimental results
We follow the experimental protocols developed in our lab to obtain the protein corona formed on top of carbon and silica NPs [41]. We first characterize the pristine NPs by Transmission Electron Microscope (TEM). By using an electron beam on dry samples, TEM provides high-resolution images and characterizes the size populations of NPs (Fig.2a and d). To study the samples in the colloidal suspension, we use Dynamic Light Scattering (DLS). In DLS, a laser tracks the movement of NPs due to Brownian diffusion, allowing the measurement of the autocorrelation function and diffusion coefficient, which in turn determine the particles’ hydrodynamic size distribution and polydispersity index (PDI) directly on the colloidal suspension. We find that both CNPs and SiO2NPs are monodisperse in aqueous solution and stable against aggregation (Fig.2a and d).
Next, we check whether the solutions of NPs interacting with Fib are stable at room temperature. In fact, it is well known that the addition of proteins in solution can induce NP aggregation [42] or disaggregation [43], and aggregation in simple media can occur at much lower protein concentrations than in complex media with many components [44].
Therefore, we use various techniques to analyze the stability of NP-protein corona complexes against aggregation after incubation at different protein-to-NP weight ratios (WRs), ranging from 0.1 to 1.0. By employing DLS and Differential Centrifugal Sedimentation (DCS), we do not observe aggregates for CNPs at the tested WRs (Fig.2 b and c). However, for the same range of WRs, aggregation systematically occurs with SiO2NP (Fig.2 e and f).
Therefore, the difference in the NPs’ chemistry likely impacts protein adsorption and aggregative propensity differently. Next, we set our experimental conditions to WR=1 and compare the cases of both CNP (colloidally stable) and SiO2NP (colloidally unstable) with a free Fib solution at various temperatures.
@@
Next, our DLS analysis indicates that the hydrodynamic diameter of free Fib and Fib with silica NPs varies with temperature. In both cases, there is an elevated PDI¿0.4 corresponding to multimodal distributions (Fig.2 e and f), with aggregation occurring at 52C and above.
In contrast, the sample containing Fib and carbon NPs shows a low PDI¡0.2, associated with an unimodal distribution, at any temperature (Fig.2 b and c). However, at 52C, it exhibits a shift in PDI that may be related to a gradual denaturation of Fib, which affects the solution’s colloidal stability.
@@ 2.1.2 UV Resonance Raman Spectroscopy
UVRR spectroscopy is a powerful tool for detecting structural changes in organic biomolecules, such as proteins and peptides, without the need for any chemical label [46]. Compared to spontaneous Raman spectroscopy using visible excitation sources, the UVRR technique offers undeniable advantages. The visible Raman profiles of proteins typically result in a complex superposition of several signals, making analysis difficult. Thanks to the selective resonance enhancement achieved by appropriately tuning the excitation wavelength in the deep UV range (200-300 nm), UVRR spectra of proteins can be simplified, and the signals arising from selected chromophores or molecular portions of the molecule can be more easily disentangled [47, 48, 49]. In this regard, using Synchrotron Radiation (SR) as an excitation source for UVRR experiments provides the added advantage of fine-tuning the excitation wavelength to select the best working conditions for the system under investigation [50, 51, 52]. Additionally, the high sensitivity of the UVRR technique is crucial for investigating proteins in more diluted conditions compared to visible Raman spectroscopy while maintaining satisfactory quality in the spectra. This, in turn, provides the opportunity to simultaneously extract information from the UVRR spectra of proteins in aqueous solutions, both from the Raman signals attributed to the polypeptide solute and the water molecules in the hydration shell [53, 51].
@@
We collect the UVRR spectra for Fib in PBS buffer, both in the presence and absence of NPs, using 226 nm as the excitation source (Fig. 4 a-c). This excitation energy falls within the absorption range of the aromatic rings of Tyrosine (Tyr) and Tryptophan (Trp) residues, resulting in UVRR spectra of Fib that are predominantly characterized by the vibrational signals of these aromatic amino acids [47, 54] (Fig. 4 a).
@@
In the Raman spectrum of the Fib protein, we can identify several key resonance Raman features associated with the Trp vibrational modes, including signals at 1012 $\mathrm{cm^{-1}}$ (Trp W16), which are attributed to the symmetric benzene/pyrrole out-of-phase breathing, at 1340-1360 $\mathrm{cm^{-1}}$ (Trp W7), which is attributed to the Fermi resonance between N1-C8 stretching in the pyrrole ring and combination bands of out-of-plane bending, and at 1554 $\mathrm{cm^{-1}}$ (Trp W3), which relates to the C2-C3 stretching mode of the pyrrole ring [55]. In the same spectral range of 1000-1800 $\mathrm{cm^{-1}}$, Tyr Raman bands are also detected, specifically the signals at 1176 $\mathrm{cm^{-1}}$ (Tyr Y9a), which are associated with in-plane C-H bending, and at 1616 $\mathrm{cm^{-1}}$ (Tyr 8a), which is due to the in-plane ring stretching of Tyr [56]. Additionally, a signal at approximately 1770 $\mathrm{cm^{-1}}$ is observed in the UVRR spectrum of Fib, likely corresponding to a C=O stretching vibration.
@@
Figure 4: UVRR spectra for free Fib, and Fib with NPs. The 226 nm-excited UVRR spectra are collected for Fib 2mg/mL in PBS pH 7.4 (a), Fib with CNPs (b), and Fib with SiO2NPs (c) as a function of temperature, reported in the two spectral ranges 1000-1800 and 2800-3800 $\mathrm{cm^{-1}}$. The main vibrational features in the spectrum of Fib are labelled in the panel (a). Temperature range and steps are indicated in the colour scale at the right. Dashed lines in panels (b) and (c) represent the spectra of Fib alone at the lowest temperature of 22C.
@@
Since the aromatic rings of Trp and Tyr can participate in hydrophobic, cation-$\mathrm{\pi}$, and hydrogen-bond interactions with nearby residues or surrounding solvent molecules [55], the spectral feature parameters of the Raman bands associated with these residues provide valuable insights into their local environment. For instance, it has been reported that the vibrations of aromatic ring side chains in proteins can serve as sensitive markers for changes occurring in the tertiary structure of the proteins during unfolding and aggregation [26, 25, 57, 58, 59, 60].
@@
It is noteworthy that, in the spectral range of 1000-1800 $\mathrm{cm^{-1}}$ (Fig. 4 a), the Raman signal from the water solvent is practically negligible compared to the vibrational signals attributed to the Fib protein. Conversely, the intense spectral distribution characteristic of the OH stretching profile of water entirely dominates the high frequency range between 2800 and 3800 $\mathrm{cm^{-1}}$ in the UVRR spectrum of Fib.
@@
Figure 5: Integrated area of some characteristic bands in the UVRR spectra in Fig. 4. (a) OH band area for Fib in PBS (black circles) and with CNPs (red pentagons), as well as SiO2NPs (green squares). (b)-(e) Areas of the peaks in the frequency intervals reported in Table 1 for Fib in PBS and with CNPs and SiO2NPs. In each panel, the frequency range is indicated as a label, and the symbols are the same as in panel (a). Where indicated, the experimental data have been fitted with a sigmoidal curve (full black lines).
The integrated area of the OH stretching band for Fib alone (Fig. 5 a) exhibits a clear sigmoidal trend as a function of temperature, indicating a transition that occurs at approximately 57C, likely associated with structural transformations for the Fib protein. The sudden increase in this Raman signal above 50C likely marks the unfolding and aggregation of the protein driven by the temperature rise. The sigmoidal trend suggests that protein denaturation begins at around 50C and rapidly increases until 70C, at which point the intensity of the OH band area stabilizes. Consequently, aggregates may be forming within this temperature range.
When CNPs are added to Fib’s solution, the OH profile significantly decreases in intensity but retains its shape (Fig. 4 b). This may suggest that water is not able to penetrate the protein corona, likely due to protein unfolding on the surface of the NP. Additionally, the corresponding OH-band integrated area shows little to no dependence on $T$ (Fig. 5 a). This finding suggests that the Fib corona is so strongly associated with particles that even with a significant increase in the energy of the solution due to heating, the proteins remain bound to the CNP. Furthermore, the analysis provides no evidence of aggregation.
When SiO2NPs are in solution with Fib, we observe that the OH-band integrated area is intermediate between the two previous cases, with a $T$-dependence that is weaker than for free Fib (Fig. 5 a). Furthermore, the increase in the area occurs at a temperature close to that of the free Fib. This behavior suggests that Fib absorbed onto SiO2NPs undergoes unfolding and aggregation similar to free Fib. Nevertheless, the analysis does not allow us to quantify the amount of unfolding and aggregation relative to the bulk case.
We also analyze the temperature dependence of the integrated area of some characteristic Raman signals in the fingerprint region 1000-1800 $\mathrm{cm^{-1}}$ assigned to Fib (Fig. 5 c-e). In particular, the unfolding and aggregation of proteins should alter the local packing of side chains and the solvent exposure of the residues, affecting the surrounding environment of Trp amino acids and the resonance Raman cross-section of the associated bands [54, 61, 62]. Therefore, by comparing their relative intensity for Fib in the absence and presence of NPs, we can qualitatively assess the effect of NPs on the $T$-dependence of the bands (Fig. 5 c-e).